OpenStudy (anonymous):

A water tank has the shape of a cone (ice cream cone). The tank is 10m high and has a radius of 3m at the top. If the water is 5m deep (in the middle) what is the surface area of the top of the water?

OpenStudy (chaise):

Area of a circle is (the surface area of the top of the cone) \[\pi \times r ^{2}\] Substitute in your values. Your r^2 value is 9 Your answer is 9xpi Surface area (approx) = 28m^2 I hope this is what you're asking? If this isn't; can you draw a picture so I can help you further?

OpenStudy (anonymous):



OpenStudy (anonymous):

that's how it look like, the top circle have 3m radius. The question is asking for the smaller circles surface area.

OpenStudy (chaise):

Okay. I see now. Because it is a cone. The circles and their radius will be DIRECTLY proportional. 10m high : 3m radius 5m high : x radius Because when we get from 10m --> 5m are dividing by two. The same is done to the radius. I hope this is right. So the radius is 1.5m, does this make sense? Failing this - I'm pretty sure you could use trigonometry and creating 'inivisble' right angle triangles. Assuming the radius is 1.5m 1.5^2 x pi = 2.25 x pi = (approx) 6.8m^2

OpenStudy (anonymous):



OpenStudy (anonymous):

there are two similar tiangles, so you can get the radius of the surface of the water by x/3=5/10, x=1.5 so S=1.5^2 pi=2.25pi